### v2

# First-Order Sparse Convex Optimization: Better Rates with Sparse Updates

Abstract
It was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution. However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving also the runtime of each iteration, which may still be prohibitively high for high-dimensional problems. In this work, we establish that linear convergence rates which depend on this improved condition number can be obtained using only sparse updates, which may result in overall significantly improved running times. Moreover, our methods are considerably easier to implement.
1 Introduction
Convex optimization techniques have been extremely successful in problems where the goal is to recover a sparse signal given certain measurements, e.g., [4, 5, 6, 7, 8, 9, 17, 18]. This work concerns certain canonical and popular convex relaxations to continuous optimization problems in which the goal is to find some optimal solution which is assumed to be sparse. We begin by presenting our two main examples.
Example 1: $\ell_{1}$-constrained minimization for entry-wise sparse recovery.
$\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~\|{{\mathbf{x}}}\|_{1}\leq R}f({\mathbf{x}}),$ (1)
[…]
Example 2: trace-constrained minimization for low-rank matrix recovery.
$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}),$ (2)
where $\mathbb{S}^{n}$ is the space of real symmetric $n\times{}n$ matrices and the relation ${\mathbf{X}}\succeq 0$ denotes ${\mathbf{X}}$ is positive semidefinite. Here we also assume $f$ is convex and continuously differentiable, and that Problem (2) indeed admits a unique optimal solution ${\mathbf{X}}^{*}$ such that $\textrm{rank}({\mathbf{X}}^{*})<<n$.
[…]
In recent works [16, 14], it was established that for convex relaxations of sparse optimization problems such as Problems (1) and (2), and under the quite plausible assumption that the optimal solution is indeed sparse/low-rank, it is possible to obtain linear convergence rates which depend on a mixed-norm condition number of the form $\frac{\beta_{1}s}{\alpha_{2}}$, where $\beta_{1}$ is the Lipchitz continuity constant of the gradient w.r.t. the $\ell_{1}$ norm, $\alpha_{2}$ is the quadratic growth constant w.r.t. the $\ell_{2}$ norm, and $s$ is an upper-bound on the sparsity of the optimal solution. This condition number could indeed be far better than the pure $\ell_{2}$ or pure $\ell_{1}$ condition numbers. For instance, in the context of Problem (1), consider the case that $f$ is convex quadratic of the following very simple form:
$f({\mathbf{x}})=\frac{1}{2}({\mathbf{x}}-{\mathbf{x}}^{*})^{\top}{\mathbf{A}}({\mathbf{x}}-{\mathbf{x}}^{*});\quad{\mathbf{A}}:=\alpha{\mathbf{I}}+\mathbf{1}\mathbf{1}^{\top},$ (3)
where $\mathbf{1}$ denotes the all-ones vector in $\mathbb{R}^{n}$ and $\alpha$ is a positive scalar.
Indeed if $\|{{\mathbf{x}}^{*}}\|_{1}\leq R$, ${\mathbf{x}}^{*}$ is the unique optimal solution to Problem (1) for this $f$. One can easily observe that the Euclidean condition number for this problem is given by $\frac{\beta_{2}}{\alpha_{2}}=\frac{\lambda_{\max}({\mathbf{A}})}{\lambda_{\min}({\mathbf{A}})}=\frac{n+\alpha}{\alpha}$. On the other-hand, a simple calculation shows that the $\ell_{1}$-Lipchitz continuity constant of the gradient is given by $\beta_{1}=\max_{i,j}|{{\mathbf{A}}_{i,j}}|=1+\alpha$. To bound the $\ell_{1}$-quadratic growth constant we observe that due the the structure of the matrix ${\mathbf{A}}$,
$\forall{\mathbf{x}}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\sum_{i=1}^{n}{\mathbf{x}}_{i}-\sum_{j=1}^{n}{\mathbf{x}}^{*}_{j}}\right)^{2}.$
In particular, assuming ${\mathbf{x}}^{*}\geq 0$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$ (i.e., ${\mathbf{x}}^{*}$ lies on the boundary of the $\ell_{1}$ ball, as often expected) we have that
$\forall{\mathbf{x}}\in\{{\mathbf{y}}\in\mathbb{R}^{n}~|~{\mathbf{y}}\geq 0,\|{{\mathbf{y}}}\|_{1}=R\}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}.$
Thus, in worst-case, the $\ell_{1}$-quadratic growth constant $\alpha_{1}$ is not better than $\frac{\alpha}{n}$, which implies that the $\ell_{1}$-condition number is $\frac{\beta_{1}}{\alpha_{1}}=\frac{n+n\alpha}{\alpha}$, which is no better than the Euclidean one.
However, in the example above, if ${\mathbf{x}}^{*}$ satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, the sparsity-based mixed-norm condition number satisfies $\frac{\beta_{1}s}{\alpha_{2}}=\frac{s(1+\alpha)}{\alpha}$, which is in particular independent of the ambient dimension, and thus may lead to far-better (worst-case) convergence rates for suitable first-order methods.
The example in Eq. (3) could be directly extended also to Problem (8) by taking $f({\mathbf{X}})=\frac{\alpha}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\textrm{Tr}({\mathbf{X}}-{\mathbf{X}}^{*})}\right)^{2}$, where $\|{\cdot}\|_{2}$ here is the Euclidean norm in $\mathbb{S}^{n}$, and instead of the entry-wise $\ell_{1}$ norm we consider the Schatten 1-norm (i.e., the $\ell_{1}$ norm w.r.t. the eigenvalues), and $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$.
The recent works [16, 14]111while [16, 14] consider a stochastic convex optimization setting in which the objective function $f$ is given as an expectation w.r.t. unknown distribution from which we can sample, here our focus is on the standard deterministic first-order convex optimization model indeed lead to convergence rates of the form $\tilde{O}\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ [16] and even an accelerated rate of the form $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$ [14] (the $\tilde{O}(\cdot)$ notation hides certain poly-logarithmic factors), by applying a restarted non-Euclidean gradient method222these methods rely on Bregman proximal gradient computations with the Bregman distance induced by a squared $\ell_{p}$ norm for an appropriate choice of $p$ that best leverages the strong convexity w.r.t. the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsify intermediate iterates, see for instance Corollary 5 in [14].
While these non-Euclidean gradient methods lead to superior convergence rates, they do not leverage the sparsity of optimal solutions towards additional computational gains, and in particular may still require prohibitive running times per iteration when the dimension is large. For instance, in the case of Problem (1), these may suffer from prohibitive running time for updating the gradient direction from one iteration to the next. In the case of Problem (2) the time required by these methods to project onto the positive semidefinite cone ($O(n^{3})$ in standard implementations) may be prohibitive.
In this work we present a new and simple approach towards obtaining linear convergence rates that depend on the improved condition number $\frac{\beta_{1}s}{\alpha_{2}}$ (though our method only obtains rates that are linear in it and not accelerated rates which depend only on its square root). The main novel aspect is that our algorithm only applies sparse updates (with the same level of sparsity assumed for the optimal solution), which often allows for significant improvements in the runtime of each iteration. In particular, our algorithm updates its iterates similarly to the classical Frank-Wolfe algorithm [15]:
${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}\quad\textrm{for some}\quad\eta\in[0,1],$ (4)
where ${\mathbf{v}}_{t}$ is a feasible point with the same level of sparsity assumed for ${\mathbf{x}}^{*}$.
For instance, in case of Problem (1) with a quadratic objective function $f({\mathbf{x}})={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, the fact that ${\mathbf{v}}_{t}$ is $s$-sparse, reduces the time to update the gradient direction from $O(n^{2})$ to $O(sn)$. In the case of Problem (2), since $\textrm{rank}({\mathbf{v}}_{t})\leq s$, computing it won’t require a full eigen-decomposition ($O(n^{3})$ time in standard implementations), but only computing the $s$ leading components in the eigen-decomposition, which is far more efficient.
Two additional advantages which make our approach considerably simpler (compared to [16, 14]) are: I. it does not require restarts, i.e., it is a single-loop algorithm, and II. the updates in our algorithm (the way ${\mathbf{v}}_{t}$ in Eq. (4) is computed) are Euclidean, which often leads to significantly simpler computations.
Compared to [16, 14], we do have an additional mild assumption, which in the case of Problems (1) and (2) above, requires that the optimal solution is on the boundary in the sense that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ or $\textrm{Tr}({\mathbf{X}}^{*})=R$, respectively. We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm regularization is redundant).
The rest of the paper is organized as follows. In Section 2 we present a unified setup and optimization problem which captures the spare convex optimization problems we are interested in, a generic algorithm for solving this problem, and its convergence analysis. In Section 3 we instantiate our generic algorithm in the special cases of Problems (1) and (2). In Section 4 we give a simple numerical demonstration for the performance of our method on Problem (1). Finally, in Section 5 we outline several open questions.
2 Unified Setup, Algorithm, and Convergence Analysis
Our formal unified setup for tackling the sparse convex optimization problems under consideration involves the following convex optimization problem:
$\min_{{\mathbf{x}}\in\mathcal{K}}f({\mathbf{x}}),$ (5)
[…]
$\forall{\mathbf{x}},{\mathbf{y}}\in\mathcal{K}:\quad\|{\nabla{}f({\mathbf{x}})-\nabla{}f({\mathbf{y}})}\|_{*}\leq\beta\|{{\mathbf{x}}-{\mathbf{y}}}\|,$
where $\|{\cdot}\|_{*}$ denotes the dual norm.
We let $\|{{\mathbf{x}}}\|_{2}=\sqrt{\langle{{\mathbf{x}},{\mathbf{x}}}\rangle}$ denote the Euclidean norm.
We shall assume Problem (5) admits a unique optimal solution ${\mathbf{x}}^{*}$, and satisfies a quadratic growth condition with parameter $\alpha_{2}>0$ w.r.t. the Euclidean norm, i.e.,
$\forall{\mathbf{x}}\in\mathcal{K}:\quad\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}\leq\frac{2}{\alpha_{2}}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right).$ (6)
We now make an additional assumption that further specializes Problem (5). In Section 3 we show that in the case of Problems (1) and (2) (and certain generalizations), Assumption 1 is satisfied once we make the plausible assumption that the norm constraint is tight for the optimal solution ${\mathbf{x}}^{*}$.
Assumption 1.
There exists a closed subset $\mathbb{E}_{s}\subset\mathbb{E}$ such ${\mathbf{x}}^{*}\in\mathcal{K}_{s}:=\mathcal{K}\cap\mathbb{E}_{s}$, and there exist scalars $s,s_{\mathcal{K}}^{*}\geq 1$ such that,
  1. 1.
    for all ${\mathbf{x}},{\mathbf{y}}\in\mathbb{E}_{s}$: $\|{{\mathbf{x}}-{\mathbf{y}}}\|\leq\sqrt{s}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}$,
  2. 2.
    for all ${\mathbf{x}}\in\mathcal{K}$: $\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|\leq\sqrt{s^{*}_{\mathcal{K}}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}$,
  3. 3.
    for all ${\mathbf{x}}\in\mathcal{K}$: $\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|=\arg\min\nolimits_{{\mathbf{y}}\in\mathcal{K}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|_{2}$ (note the argmin need not be a singleton),
In the context of Problems (1) and (2) we should think of $\mathbb{E}_{s}$ as the restriction of the space $\mathbb{E}$ to vectors which are $s$-sparse (in some sense), and accordingly $\mathcal{K}_{s}$ is the restriction of $\mathcal{K}$ to $s$-sparse vectors. Additionally, while not part of Assumption 1, we shall implicitly assume that it is computationally efficient to compute a Euclidean projection onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$ (in case the projection is not a singleton, compute one element in the set of possible projections). Accordingly, we denote by $\Pi_{\mathbb{E}_{s}}[\cdot]$ and $\Pi_{\mathcal{K}_{s}}[\cdot]$ the projection operation onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$, respectively.
We can now present our algorithm for solving the generic Problem (5) (under the assumptions listed above). Our algorithm is composed of three simple ingredients. First, we use a Frank-Wolfe like scheme [15] in which the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ are updated by taking a convex combination between the current feasible iterate and a new feasible point, as already shown in Eq. (4). Second, and differently from the classical Frank-Wolfe method in which the new point ${\mathbf{v}}_{t}$ is the output of a linear optimization oracle, here we take it to be a Euclidean proximal gradient step w.r.t. to restricted set $\mathcal{K}_{s}$ (instead of w.r.t. the original set $\mathcal{K}$). This is similar to approach taken in previous works [1, 10, 12, 13] (also referred to as a weak proximal oracle [10, 13]). Third, the $\ell_{2}$-term in this proximal gradient step is not w.r.t. to the current iterate ${\mathbf{x}}_{t}$, but with respect to its projection onto $\mathbb{E}_{s}$, which should be understood as a hard-thresholding (or sparsification) of ${\mathbf{x}}_{t}$.
Algorithm 1 Frank-Wolfe with sparse proximal oracle and hard-thresholding
${\mathbf{x}}_{1}\leftarrow$ some arbitrary point in $\mathcal{K}$
for $t=1,2\dots$ do
  $\widehat{{\mathbf{x}}}_{t}\leftarrow\Pi_{\mathbb{E}_{s}}[{\mathbf{x}}_{t}]$
  ${\mathbf{v}}_{t}\leftarrow\arg\min\nolimits_{{\mathbf{v}}\in\mathcal{K}_{s}}\langle{{\mathbf{v}}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+s\beta\eta\|{{\mathbf{v}}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}$equivalently, ${\mathbf{v}}_{t}\leftarrow\Pi_{\mathcal{K}_{s}}[\widehat{{\mathbf{x}}}_{t}-\frac{1}{2s\beta\eta}\nabla{}f({\mathbf{x}}_{t})]$}
  ${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}$
end for
Theorem 1.
Consider the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ of Algorithm 1 with step-size $\eta=\frac{\alpha_{2}}{4\beta(4s+s^{*}_{\mathcal{K}})}$. Then,
$\forall t\geq 1:\quad f({\mathbf{x}}_{t+1})-f^{*}\leq\left({f({\mathbf{x}}_{t})-f^{*}}\right)\left({1-\frac{\alpha_{2}}{8\beta(4s+s^{*}_{\mathcal{K}})}}\right).$
Proof.
Fix some iteration $t$ of the algorithm. The Lipchitz continuity of the gradient implies the following well known inequality:
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\frac{\beta\eta^{2}}{2}\|{{\mathbf{v}}_{t}-{\mathbf{x}}_{t}}\|^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Since $\widehat{{\mathbf{x}}}_{t}$ and ${\mathbf{v}}_{t}$ are in $\mathbb{E}_{s}$, using Item 1 in Assumption 1 we have that,
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Using the definition of ${\mathbf{v}}_{t}$ in the algorithm and the assumption that ${\mathbf{x}}^{*}\in\mathcal{K}_{s}$ we have that,
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Using the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm, Item 3 and then Item 2 in Assumption 1 we have that,
$\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}\leq\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|^{2}\leq s_{\mathcal{K}}^{*}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}.$
Combining the last two inequalities we have that,
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}s^{*}_{\mathcal{K}}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(2s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}+2\beta\eta^{2}s\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2},$
where the last inequality follows from the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm.
Using both the convexity and quadratic growth of $f$ (Eq. (6)), we have that
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})-\eta\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right)+\frac{2\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})}{\alpha_{2}}\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right).$
Subtracting $f({\mathbf{x}}^{*})$ from both sides and plugging-in our choice of step-size $\eta$, completes the proof.
3 Applications
3.1 $\ell_{1}$-ball setup
We recall the first problem in our Introduction:
$\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~\|{{\mathbf{x}}}\|_{1}\leq R}f({\mathbf{x}}),$ (7)
[…]
We shall require the following simple lemma (a proof is brought in the appendix for completeness).
Lemma 1.
Let ${\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{n}$ such that $\|{{\mathbf{y}}}\|_{1}\leq\|{{\mathbf{x}}}\|_{1}$. Then,
$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}\leq 2\sqrt{\textrm{nnz}({\mathbf{x}})}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
Let $s$ be such that $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\textrm{nnz}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\|{{\mathbf{x}}}\|_{1}\leq R,~\textrm{nnz}({\mathbf{x}})\leq s\}$. In light of Lemma 1 and the assumption that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=2s$.
Note that computing the hard-thresholding step in Algorithm 1 (the point $\widehat{{\mathbf{x}}}_{t}$) is straightforward by zeroing-out all but the largest (in absolute value) entries in ${\mathbf{x}}_{t}$.
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, by computing the Euclidean projection of the point ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$ onto the $s$-sparsified $\ell_{1}$-ball $\mathcal{K}_{s}$, is done by computing the Euclidean projection of the vector in $\mathbb{R}^{s}$ which corresponds to the $s$-largest (in absolute value) entries in ${\mathbf{z}}_{t}$ onto the $s$-dimensional $\ell_{1}$-ball of radius $R$, and then appending zeros (in the entries which correspond to the $n-s$ lower entries in ${\mathbf{z}}_{t}$) to obtain again a point the $n$-dimensional $\ell_{1}$-ball of radius $R$. See [3] for proof of the correctness of this procedure.
Overall, aside from the gradient computation, each iteration of Algorithm 1 requires $O(n\log{}s)$ time (which is the time to find the $s$ largest entries in an array of length $n$ — the most time consuming operation).
Corollary 1.
When applied to Problem (7) with an optimal solution ${\mathbf{x}}^{*}$ which satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside of gradient computations, the runtime of each iteration is $O(n\log{s})$. In particular, if $f$ is (convex) quadratic, the overall runtime of each iteration is $O(sn)$.
Remark 1.
It is straightforward to extend the result above to the case in which ${\mathbf{x}}$ is constrained to be in $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{x}}~|~\|{{\mathbf{x}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{n}$ is closed and convex, as long as ${\mathbf{x}}^{*}$ satisfies $\|{{\mathbf{x}}}\|_{1}=R$. For instance when $\mathcal{C}=\{{\mathbf{x}}~|~{\mathbf{x}}\geq 0,\sum_{i=1}^{n}{\mathbf{x}}_{i}=R\}$, we have that $\mathcal{K}$ is simply a scaled simplex for which the assumption $\|{{\mathbf{x}}^{*}}\|_{1}=R$ always holds. Moreover, when $\mathcal{C}$ is closed under permutation of coordinates (e.g., an $\ell_{p}$ ball or its non-negative restriction), the sparse proximal step (computation of ${\mathbf{v}}_{t}$ in Algorithm 1) admits efficient implementation as detailed in [3].
3.2 Bounded positive semidefinite cone setup
We recall the second problem in our Introduction:
$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}).$ (8)
[…]
The following lemma is analogues to Lemma 1. A proof is given in the appendix for completeness.
Lemma 2.
Let ${\mathbf{X}},{\mathbf{Y}}\in\{{\mathbf{Z}}\in\mathbb{S}^{n}~|~{\mathbf{Z}}\succeq 0\}$ such that ${\mathbf{X}}\neq\mathbf{0}$, $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$. Then,
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq 4\sqrt{\textrm{rank}({\mathbf{X}})}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$
Let $s$ be such that $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~\textrm{rank}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\|{{\mathbf{X}}}\|_{1}\leq R,~\textrm{rank}({\mathbf{X}})\leq s\}$. In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=16s$.
[…]
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, is done as follows: compute the $s$ top components (corresponding to the largest signed eigenvalues) in the eigen-decomposition of the matrix ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$, and denote it as $\sum_{i=1}^{s}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, ${\mathbf{u}}_{i}\in\mathbb{R}^{n},i=1,\dots,s$. Then, project the vector of eigenvalues $(\lambda_{1},\dots,\lambda_{s})$ onto the $s$-dimensional downward closed simplex $\mathcal{S}_{R}:=\{{\mathbf{y}}\in\mathbb{R}^{s}~|~{\mathbf{y}}\geq 0,~\sum_{i=1}^{s}{\mathbf{y}}_{i}\leq R\}$ to obtain the new vector of eigenvalues $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$. Then, we set ${\mathbf{v}}_{t}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ (i.e., the eigenvectors remain unchanged). For the correctness of this procedure see Lemma 3 in the appendix.
[…]
Corollary 2.
[…]
Remark 2.
As in Remark 1, here also we can extend Problem (8) to include slightly more involved constraints of the form $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R\}$, where $\mathcal{C}\subset\mathbb{S}^{n}$ is closed and convex, as long as $\textrm{Tr}({\mathbf{X}}^{*})=R$. In particular, when $\mathcal{C}$ is a ball w.r.t. a Schatten $p$-norm ($\ell_{p}$ norm w.r.t. the eigenvalues), one can extend Lemma 3 accordingly.
4 Some Numerical Evidence
We present simple numerical simulations to demonstrate the benefit of the improved convergence rate of our method, which depends on a mixed-norm condition number, in the context of the $\ell_{1}$-constrained Problem (1). Similarly to the example in Eq. (3), we set $f({\mathbf{x}})=\frac{1}{2}({\mathbf{x}}-{\mathbf{x}}^{*})^{\top}({\mathbf{I}}+3\cdot\mathbf{1}\mathbf{1}^{\top})({\mathbf{x}}-{\mathbf{x}}^{*})$, where $\mathbf{1}$ denotes the all-ones vector. For this problem we have that the $\ell_{2}$-quadratic growth (in fact, strong convexity) constant is $\alpha_{2}=1$, and the $\ell_{2}$ and $\ell_{1}$ Lipchitz continuity constants of the gradient are $\beta_{2}=1+3n$ and $\beta_{1}=4$, respectively. We compare the convergence rate of our Algorithm 1 to a state-of-the-art Euclidean accelerated gradient method for smooth and strongly convex optimization known as V-FISTA [2], which has worst-case convergence rate $O((1-\sqrt{\alpha_{2}/\beta_{2}})^{t})$ (see Theorem 10.42 in [2]).
We set the dimension $n=3000$, the radius of the $\ell_{1}$ ball to $R=10$. The optimal solution ${\mathbf{x}}^{*}$ is drawn at random so that each non-zero entry (the non-zero locations are also randomized) is set to ±$\pm\frac{R}{\textrm{nnz}({\mathbf{x}}^{*})}$ with equal probability, where $\textrm{nnz}({\mathbf{x}}^{*})$ is set in advanced, and for each level of sparsity we run 10 i.i.d. experiments (where in each one we resample ${\mathbf{x}}^{*}$). Accordingly, the figure below presents the average approximation error over 10 i.i.d. experiments.
When implementing our Algorithm 1, we assumed the exact sparsity of ${\mathbf{x}}^{*}$ is known, i.e., we set $s=\textrm{nnz}({\mathbf{x}}^{*})$. Additionally, observing that the theoretical choice of step-size corresponding to Theorem 1, $\eta=\frac{\alpha_{2}}{6\beta_{1}{}s}$ (see Section 3.1) is too conservative in practice, we used a larger step-size $\tilde{\eta}=\frac{\alpha_{2}}{\beta_{1}{}s}$, when computing the point ${\mathbf{v}}_{t}$ on each iteration $t$. Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in (0,1)(0,1)( 0 , 1 )). For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2].
From Figure 1, we can clearly see that Algorithm 1 is significantly faster for highly-sparse instances, strongly outperforming even a state-of-the-art accelerated method, while as ${\mathbf{x}}^{*}$ becomes less-sparse, this advantage vanishes.
(a) $\textrm{nnz}({\mathbf{x}}^{*})=10$
(b) $\textrm{nnz}({\mathbf{x}}^{*})=30$
(c) $\textrm{nnz}({\mathbf{x}}^{*})=50$
Figure 1: Comparison of Algorithm 1 and the V-FISTA algorithm for Problem (1).
5 Open Questions
We conclude with two open questions which we believe are of interest. First, it is interesting if our approach could be extended to also handle norm-regulairzed problems instead of norm-constrained problems. Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ while maintaining the sparse updates.
Acknowledgments
This work was funded by the European Union (ERC, ProFreeOpt, 101170791). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
We would like to thank Amir Beck and Marc Teboulle for fruitful discussions during the early stages of this work.
Appendix A Proof of Lemma 1
Proof.
Denote $k=\textrm{nnz}({\mathbf{x}})$. Let $S\subseteq[n]$ be the set of indices for non-zero entries in ${\mathbf{x}}$. Let ${\mathbf{x}}_{1},{\mathbf{y}}_{1}\in\mathbb{R}^{k}$ be the restrictions of ${\mathbf{x}}$ and ${\mathbf{y}}$ to the entries indexed by $S$, respectively. Accordingly, let ${\mathbf{x}}_{2},{\mathbf{y}}_{2}\in\mathbb{R}^{n-k}$ be the restrictions of ${\mathbf{x}}$ and ${\mathbf{y}}$ to the entries not indexed by $S$. It holds that,
$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}$ $=\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{y}}_{2}}\|_{1}$
$\leq\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{y}}_{2}}\|_{1}$
$=\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{y}}}\|_{1}-\|{{\mathbf{y}}_{1}}\|_{1}$
$\leq\ \|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{x}}_{1}}\|_{1}-\|{{\mathbf{y}}_{1}}\|_{1}$
$\leq 2\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}$
$\leq 2\sqrt{k}\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{2}$
$\leq 2\sqrt{k}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
Appendix B Proof of Lemma 2
Proof.
Denote $k=\textrm{rank}({\mathbf{X}})$. Let ${\mathbf{P}}\in\mathbb{S}^{n}$ denote the projection matrix onto the subspace $\textrm{Im}({\mathbf{X}})$ and let ${\mathbf{P}}_{\perp}\in\mathbb{S}^{n}$ denote the projection matrix onto the orthogonal complement of $\textrm{Im}({\mathbf{X}})$. Throughout the proof we let $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm (sum of singular values) and the Euclidean (Frobenius) norm in some space of real (not necessarily symmetric) matrices, respectively. We begin by recalling two important facts. First, for any real (not necessarily symmetric) matrix ${\mathbf{A}}$ it holds that
$\|{{\mathbf{A}}}\|_{1}\leq\sqrt{\textrm{rank}({\mathbf{A}})}\|{{\mathbf{A}}}\|_{2}.$ (9)
Second, for any two real (not necessarily symmetric) matrices of appropriate dimension (such that the multiplication is defined) ${\mathbf{A}},{\mathbf{B}}$ it holds that
$\|{{\mathbf{A}}{\mathbf{B}}}\|_{2}\leq\|{{\mathbf{A}}}\|_{\infty}\|{{\mathbf{B}}}\|_{2},$ (10)
where we let $\|{{\mathbf{A}}}\|_{\infty}$ denote the largest singular value of ${\mathbf{A}}$.
Let us denote ${\mathbf{D}}={\mathbf{X}}-{\mathbf{Y}}$. It holds that,
$\|{{\mathbf{D}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp}){\mathbf{D}}({\mathbf{P}}+{\mathbf{P}}_{\perp})}\|_{1}$
$\leq\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}+2\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}.$ (11)
We continue by upper-bounding each of the three terms in the RHS separately.
Note that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}})\leq k$. Thus, using Eq. (9) and then Eq. (10), and noting that $\|{{\mathbf{P}}}\|_{\infty}=1$, we have that
$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (12)
To bound the second term in the RHS of (B) we have,
$\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}$ $=\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}}\|_{1}\underset{(a)}{=}\textrm{Tr}({\mathbf{P}}_{\perp}{\mathbf{Y}})=\textrm{Tr}(({\mathbf{I}}-{\mathbf{P}}){\mathbf{Y}})$
$\underset{(b)}{\leq}\textrm{Tr}({\mathbf{X}})-\textrm{Tr}({\mathbf{P}}{\mathbf{Y}})=\textrm{Tr}({\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}})\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}}}\|_{1}$
$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$ (13)
where (a) holds since ${\mathbf{Y}}$ is positive semidefinite, and thus so is ${\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}$, (b) holds by the assumption that $\textrm{Tr}({\mathbf{Y}})=\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}=\textrm{Tr}({\mathbf{X}})$, and (c) is due to Eq. (12)
Finally, for the last term in the RHS of (B), noting that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp})\leq k$, using Eq. (9) and then Eq. (10) (and noting again that $\|{{\mathbf{P}}}\|_{\infty}=\|{{\mathbf{P}}_{\perp}}\|_{\infty}=1$), we have that
$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (14)
The lemma follows from plugging-in (12), (B), and (14) into the RHS of (B). ∎
Appendix C Sparse Projections onto the Bounded Positive Semidefinite Cone
Lemma 3.
Let ${\mathbf{Z}}\in\mathbb{S}^{n}$ with eigen-decomposition ${\mathbf{Z}}=\sum_{i=1}^{n}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ ($\lambda_{1}\geq\lambda_{2}\geq...$) and $s\in[n]$. A Euclidean projection of ${\mathbf{X}}$ onto $\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R,\textrm{rank}({\mathbf{X}})\leq s\}$ is given by the matrix ${\mathbf{Z}}^{\prime}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, where $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of the vector $(\lambda_{1},\dots,\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
Proof.
[…]
$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}$ $=\|{{\mathbf{Y}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}}\|_{2}^{2}-2\langle{{\mathbf{Y}}-{\mathbf{Z}}^{\prime},{\mathbf{Z}}}\rangle$
$=\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\langle{{\mathbf{Y}},{\mathbf{Z}}}\rangle+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i}$
$\geq\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\sum_{i=1}^{s}\gamma_{i}\lambda_{i}+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i},$
where the last inequality is due to von Neumann’s trace inequality.
Rearranging, we have that
$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}\geq\sum_{i=1}^{s}(\gamma_{i}-\lambda_{i})^{2}-\sum_{j=1}^{s}(\lambda_{j}^{\prime}-\lambda_{j})^{2}\geq 0,$
where the last inequality follows since $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of $(\lambda_{1},\dots\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$. ∎